OpenStudy (anonymous):

A car 250 miles due east of point A is traveling west at a constant rate of 60 miles per hour whiles another car 314 miles due west of point A is traveling east at a constant rate of 55 miles per hour. After about how many hours will the 2 cars meet? 3.1 4.9 5.4 7.6 9.7

OpenStudy (anonymous):

x:distance we have v=dx/dt so x=60t-250 and x=-55t+314 t=4.9